# PINN-MEP: Continuous Neural Representations for Minimum-Energy Path Discovery in Molecular Systems

### v4

@@ 4 BPTI Conformational Change Pathway
Bovine Pancreatic Trypsin Inhibitor (BPTI) is a clinically significant protein used as an anticoagulant in medical procedures. BPTI is 58 residues long and has complex conformational changes involving disulfide bond rearrangements paired with backbone changes.
@@
To investigate BPTI’s conformational transitions, we use as reference structures five key states from a landmark molecular dynamics (MD) simulation performed by D.E. Shaw Research on their Anton supercomputer (Shaw et al., 2010). These publicly available snapshots provide rare benchmark structures for validating our approach to BPTI’s transition mechanisms. Full details of our BPTI system preparation, explicit solvent shell construction, and hybrid neural network architecture for handling water molecules are provided in Appendix B.2.
@@
We generate multiple transition MEPs connecting two BPTI states from the D.E. Shaw simulation. The intermediate snapshots from the D.E. Shaw simulation serve as validation points; they were not used as inputs but represent physically realized conformations from the whole trajectory. To analyze this path, we compute two collective variables matching those used in the original D.E. Shaw study: (1) the disulfide torsion angle between Cys14 and Cys38, calculated using the dihedral angle formed by CB14–SG14–SG38–CB38 atoms, and (2) the backbone RMSD of residues 4-54 after mean-centering. Remarkably, our optimized path passes through all intermediate snapshots, indicating that it successfully discovers the same transition mechanism as the computationally intensive explicit-solvent MD simulation.
@@
We matched their simulation conditions by incorporating explicit water molecules in our model to ensure a meaningful comparison with the Anton simulation. We generated water shells extending 14Å from any protein atom for both the initial and final BPTI conformations. To handle water molecules during transitions, we pair proximal water molecules between start and end states, then model their movement using simple parameterized splines, rather than directly incorporating all water coordinates into the neural network. This hybrid neural-spline approach allows us to include explicit solvent effects while keeping computational costs low. The specific details of our water shell construction and matching algorithm are described in Appendix B.2.
The system was modeled using the AMBER99sb-ILDN force field (Lindorff-Larsen et al., 2010) with the TIP3P water model (Jorgensen et al., 1983; Neria et al., 1996). Each minimum-energy path was computed in approximately 15 minutes on a single A6000 GPU, compared to the D.E. Shaw simulation, which required approximately 1.3 million node-hours on their specialized Anton supercomputer to generate the 1.03 millisecond trajectory capturing the same conformational changes. While our approach does not provide the same statistical ensemble information as equilibrium MD, it efficiently captures the key transition pathway. It’s worth noting that even the reference millisecond-scale simulation may not have fully achieved equilibrium sampling of the conformational space. Appendix B.2.7 provides a detailed computational performance comparison.
@@ 5 Limitations
While our feedforward neural architecture can model conformational transitions, it could benefit from incorporating domain-specific structural priors to improve training efficiency and convergence speed. Furthermore, physical force fields introduce challenges through high-energy spikes that can cause training instability and divergence in some runs. Our gradient clipping approach mitigates this issue but doesn’t universally prevent training failures.
The method has been validated primarily on small to medium-sized protein systems, and scaling to larger multi-domain proteins could require additional optimization strategies. Our approach is also constrained to systems with differentiable force fields, but this limitation diminishes as more implementations become available.
Finally, while computationally efficient, our method focuses on identifying minimum-energy paths rather than sampling the full statistical ensemble of transition pathways, representing a trade-off between efficiency and exhaustively sampling a system.
@@ 6 Conclusion
We introduced a novel approach to molecular transition path discovery that leverages physics-informed neural networks and differentiable force fields to enable the continuous representation of conformational changes. By reformulating minimum-energy path discovery as a neural optimization problem, we demonstrated the method’s effectiveness on both the small AIB9 peptide and the explicitly solvated BPTI protein, showing its ability to discover physically meaningful transition pathways across different complexity scales.
Our approach achieves remarkable computational efficiency, reducing computational requirements by approximately six orders of magnitude compared to traditional MD simulations for the BPTI system. While our method yields a single pathway rather than the complete ensemble of transition trajectories, this represents a valuable trade-off that makes previously intractable conformational transition studies accessible on standard computational resources.
While our current implementation shows promising results, several directions for improvement remain. Future work should explore neural architectures that better respect the structural organization of protein systems. Ramakrishnan et al. (2025) investigates non-uniform sampling schemes along the path, among other innovative design decisions that might further aid convergence. Integrating pretrained generative models suggested by Raja et al. (2025) could provide better initial transition guesses, potentially accelerating convergence by starting optimization from more plausible conformational pathways. Extending the method to handle larger multi-domain proteins is also a promising frontier. Future work could also explore methods to estimate free energies and reaction rates from our discovered paths.
@@ A.1 Extended derivation of the energy-based loss function
For a minimum-energy path, the displacement tends to align either parallel or anti-parallel with the gradient direction (i.e., for path segments)
@@
A.1.1 From Onsager-Machlup to a geometric line integral
@@
with equality when is parallel or anti-parallel to , with the direction determined by whether the path is ascending or descending the energy landscape. This is precisely the defining characteristic of the minimum-energy path (MEP) - a path that follows the potential energy landscape’s gradient while avoiding high-energy regions.
@@ A.1.4 Neural network implementation
For practical optimization, we parametrize the path using a neural network with parameters . To implement our approach, we sample points along the path and minimize the energy at these sampled locations. This represents a transition from the theoretical derivation using discrete segments to a practical implementation using sampled points for more efficient optimization. We employ stochastic sampling of the progress variable to construct our loss function:
@@
By minimizing through gradient-based optimization, the neural network learns to represent low-energy pathways connecting the stable states, effectively discovering the minimum-energy path predicted by the Onsager-Machlup framework. The resulting path follows the potential energy landscape’s gradient while avoiding high-energy regions, representing the most probable transition mechanism in the overdamped limit.
A.2 Connection between Doob’s Lagrangian and Minimum-Energy Path Optimization
@@ A.2.1 Comparison with Doob’s Lagrangian Method
We provide a direct comparison between our PINN-MEP approach and the method of Du et al. in terms of computational efficiency, as evaluated by the number of force-field/energy function calls, and the maximum energies observed in the paths, corresponding to the peak of the potential barrier. To provide a rigorous comparison, we conducted benchmarks using experimental conditions identical to those reported by Du et al. for alanine dipeptide: AMBER14 force field, 300K temperature. Table 1 presents these results alongside those reported in Du et al. (2024) and their recent follow-up work Lee et al. (2025). The only exception is that we do not evaluate trajectory probability, as the assumption of optimization over a non-physical path length makes the calculation of this quantity impossible.
@@ Appendix B Comparison with Other Related Methods
While our primary benchmark comparison focuses on the Doob’s Lagrangian method due to similarity in task and system setup, examining how our approach relates to other methods for transition path discovery/minimum energy path generation is also informative. Here we present results from two related approaches in the literature, though we emphasize that direct numerical comparisons should be treated with caution due to differences in force fields, optimization objectives, and computational frameworks.
| System | Method | Total energy evals | TS energy (eV) |
| Alanine dipeptide | NEB | 8063 | 11.0014 |
@@
| Method | CVs | # FF Evals / Path | Runtime / Path |
@@
Table 2 shows results from Ramakrishnan et al. (2025) comparing Nudged Elastic Band (NEB) with their Implicit Neural Representation (INR) approach for minimum-energy path discovery. Both our PINN-MEP approach and their INR method utilize lightweight, preexisting molecular dynamics force fields, though we employ different such force fields. Their approach requires a more complex loss function that evaluates both energy and forces along with path parameterization and its gradient (path velocity). In contrast, Table 3 presents results from Raja et al. (2025), who take a fundamentally different approach by repurposing pre-trained generative models to perform sampling of transition paths. A key advantage of their method is that the generative model provides significantly better initial guesses for transition paths based on learned molecular distributions. However, their approach requires evaluating both the pretrained model (as a surrogate force field) and its gradient for their loss function. Our PINN-MEP method stands out for its simplicity, using only the energy potential in the loss function without requiring additional gradient terms, while maintaining comparable computational efficiency. These tables corroborate our findings in the Doob’s Lagrangian comparison: methods optimizing for a single minimum-energy path consistently require orders of magnitude fewer force field evaluations than approaches that characterize the full transition path ensemble.
@@ B.1 Ablation Studies
This section presents a series of ablation studies examining the impact of various components of our PINN-MEP approach. These experiments help validate our design choices and provide insights into the importance of different architectural and training elements. We conduct these ablation studies on the AIB9 system because Alanine dipeptide, as a benchmark system, is too simple to provide meaningful insights into the performance of our method. For each ablation study, we report the lowest maximum energy of the path, the mean and standard deviation of the energy, and the mean number of force field evaluations to reach the lowest maximum energy. We run each setting in the ablations five times for a maximum of 150000 batches, with a batch size of 1 and 32 frames per batch. For experiments without a spring constant between adjacent sampled frames, the batch size and frames are treated identically. Due to the nature of physical force fields, slight deviations in the path can lead to large changes in energy, leading to similar settings having very different energies. This can limit the ability to compare the performance of different settings. However, we find that the results are still directionally informative.
B.1.1 Embedding Strategy
We compared sinusoidal time embeddings with different maximum periods against simple linear projections of the progress variable. Table 4 shows sinusoidal embeddings with a period of generally outperformed other configurations.
| Embedding Type | Max Period | Mean Energy | Best Energy | Evaluations |
| Sinusoidal | 5832.29 7006.09 | -33.80 | 142400 | |
| Sinusoidal | 4939.20 2836.56 | 1388.48 | 122100 | |
| Sinusoidal | 9508.61 7297.49 | 1273.18 | 106900 | |
| Sinusoidal | 5851.25 4618.35 | 819.25 | 129600 | |
| Linear (None) | – | 9567.01 7578.20 | 303.63 | 127600 |
B.1.2 Spring Regularization
We tested the effect of chain-of-states-like spring constant regularization, as shown in Table 5. Interestingly, we found that a small amount of spring regularization () provided the best results, suggesting that a minimal amount of explicit path continuity enforcement can be beneficial when combined with our neural representation, however, this is not a requirement for the method to work and the additional hyperparameter does not necessarily be worth including.
| Spring Constant | Mean Energy | Best Energy | Evaluations |
| 368.78 416.68 | -278.09 | 130300 | |
| 8862.90 9138.59 | -155.85 | 120600 | |
| 0.0 | 8200.08 15576.30 | 240.85 | 116500 |
| 1156.70 428.50 | 791.59 | 120400 | |
| 7440.13 3339.87 | 4657.28 | 92300 | |
| 45567.30 5015.78 | 40609.10 | 64700 |
B.1.3 Network Architecture
@@
We investigated the effect of network depth and width on transition path quality. Table 6 shows that an 8-layer network with a width factor of 4 achieved the best balance between expressivity and computational efficiency, with wider networks also performing well.
@@
| Depth | Width Factor | Mean Energy | Best Energy | Evaluations |
| 8 | 4 | 2746.08 5381.57 | -168.17 | 112300 |
| 8 | 6 | 6337.10 6284.13 | -278.09 | 79300 |
| 8 | 2 | 10396.20 14039.40 | -278.09 | 113500 |
| 6 | 6 | 6189.19 7667.11 | 334.22 | 112800 |
| 6 | 4 | 9371.15 15277.50 | 240.85 | 111800 |
| 6 | 2 | 9872.01 14229.30 | 152.36 | 122500 |
| 4 | 6 | 13242.00 8432.26 | 6013.58 | 94000 |
| 4 | 4 | 16300.90 5768.37 | 8295.52 | 132100 |
| 4 | 2 | 24916.70 14933.30 | 12697.60 | 78700 |
B.1.4 AIB9 Network Architecture

### v5

@@ 4 BPTI Conformational Change Pathway
Bovine Pancreatic Trypsin Inhibitor (BPTI) is a clinically significant protein used as an anticoagulant in medical procedures. BPTI is 58 residues long (892 atoms) and has complex conformational changes involving disulfide bond rearrangements paired with backbone changes.
@@
To investigate BPTI’s conformational transitions, we use as reference structures five key states from a landmark molecular dynamics (MD) simulation performed by D.E. Shaw Research on their Anton supercomputer (Shaw et al., 2010). These publicly available snapshots provide rare benchmark structures for validating our approach to BPTI’s transition mechanisms, making it an excellent larger system to test the method. Full details of our BPTI system preparation, explicit solvent shell construction, and hybrid neural network architecture for handling water molecules are provided in Appendix B.2.
@@
We generate multiple transition MEPs connecting two BPTI states from the D.E. Shaw simulation. The intermediate snapshots from the D.E. Shaw simulation serve as validation points; they were not used as inputs but represent physically realized conformations from the whole trajectory. To analyze this path, we compute two collective variables matching those used in the original D.E. Shaw study: (1) the disulfide torsion angle between Cys14 and Cys38, calculated using the dihedral angle formed by CB14–SG14–SG38–CB38 atoms, and (2) the backbone RMSD, root mean square displacement, of residues 4-54 after mean-centering. Remarkably, as shown in Figure 3, our optimized path passes through all intermediate snapshots, indicating that it successfully discovers the same transition mechanism as the computationally intensive MD simulation.
@@
We matched their simulation conditions by incorporating explicit water molecules in our model to ensure a meaningful comparison with the Anton simulation. We generated water shells extending 14Å from any protein atom for both the initial and final BPTI conformations. This increases the number of atoms from 892 to 8,300. To handle water molecules during transitions, we pair proximal water molecules between start and end states using the Hungarian algorithm (Kuhn, 1955). We then model their movement using simple parameterized splines, rather than directly incorporating all water coordinates into the neural network. This hybrid neural-spline approach allows us to include explicit solvent effects while keeping computational costs low. The specific details of our water shell construction and matching algorithm are described in Appendix B.2.
The system was modeled using the AMBER99sb-ILDN force field (Lindorff-Larsen et al., 2010) with the TIP3P water model (Jorgensen et al., 1983; Neria et al., 1996). Each MEP was computed in approximately 15 minutes on a single A6000 GPU, requiring only force field evaluations compared to the D.E. Shaw simulation, which conducted approximately 412 billion force field evaluations to generate the 1.03 millisecond trajectory capturing the same conformational changes. This represents a six orders of magnitude reduction in computational effort.
While our approach does not provide the same statistical ensemble information as equilibrium MD, it efficiently captures the key transition pathway, which is often the primary mechanistic insight sought by MD practitioners. For most scientifically interesting systems—typically large biomolecular complexes—equilibrium sampling of the complete conformational space or transition path ensemble remains computationally prohibitive, making identification of dominant pathways particularly valuable. It’s worth noting that even the reference millisecond-scale simulation may not have fully achieved equilibrium sampling of the conformational space. Appendix B.2.7 provides a detailed computational performance comparison.
@@ 5 Limitations
While our feedforward neural architecture can model conformational transitions, it could benefit from incorporating domain-specific structural priors to improve training efficiency and convergence speed. Furthermore, physical force fields introduce challenges through high-energy spikes that can cause training instability and divergence in some runs. Our gradient clipping approach, outlined in B.2.5, largely mitigates this issue but doesn’t universally prevent training failures.
The method has been validated primarily on small to medium-sized protein systems, and scaling to larger multi-domain proteins could require additional optimization strategies. Our approach is also constrained to systems with suitable differentiable force fields, but this limitation diminishes as more implementations become available.
Finally, while computationally efficient, our method focuses on identifying MEPs rather than sampling the full statistical ensemble of transition pathways, representing a trade-off between efficiency and exhaustively sampling a system. However, as noted in the BPTI section 4, this mechanistic insight is often a key objective for MD practitioners studying complex biomolecular transitions.
@@ 6 Conclusion
We demonstrated that reformulating MEP discovery as a continuous neural optimization problem enables efficient discovery of molecular transition mechanisms with unprecedented computational efficiency. Our approach transforms the complex Onsager-Machlup action functional into a straightforward energy-based loss function that eliminates the need for force and path gradient computations required by related methods, which are known to induce optimization challenges in PINNs. This work provides a practical tool for molecular dynamics practitioners that captures essential mechanistic insights without requiring specialized hardware or extensive simulation time.
While our current implementation shows promising results, several concrete directions for improvement remain. Future work should explore neural architectures that incorporate domain-specific structural priors beyond simple MLPs with sinusoidal/Fourier embeddings. Related work by Ramakrishnan et al. (2025) investigates non-uniform sampling schemes along the path that might further aid convergence for paths that constitute dramatic rearrangements. Integrating pretrained generative models suggested by Raja et al. (2025) could provide better initial transition guesses, potentially accelerating convergence by starting optimization from conformations that already respect molecular geometry. Extending the method to handle larger multi-domain proteins is also a frontier of great scientific interest, which might bring new challenges. Future work could also explore ways to estimate free energies and reaction rates from our discovered paths.
@@ A.1 Extended Derivation of the energy-based Loss Function
For a MEP, the displacement tends to align either parallel or anti-parallel with the gradient direction (i.e., for path segments)
@@
A.1.1 From the Onsager-Machlup Action to a geometric Line Integral
@@
with equality when is parallel or anti-parallel to , with the direction determined by whether the path is ascending or descending the energy landscape. This is precisely the defining characteristic of the MEP - a path that follows the potential energy landscape’s gradient while avoiding high-energy regions.
@@ A.1.4 Neural Network Implementation
For practical optimization, we parametrize the path using a neural network with parameters . To implement our approach, we sample points along the path and minimize the energy at these sampled locations. This represents a transition from the theoretical derivation using discrete segments to a practical implementation using sampled points for more efficient optimization. We employ stochastic sampling of the progress coordinate to construct our loss function:
@@
By minimizing through gradient-based optimization, the neural network learns to represent low-energy pathways connecting the stable states, effectively discovering the MEP predicted by the Onsager-Machlup framework. The resulting path follows the potential energy landscape’s gradient while avoiding high-energy regions, representing the most probable transition mechanism in the overdamped limit.
A.2 Connection between Doob’s Lagrangian and MEP Optimization
@@ A.2.1 Comparison with Doob’s Lagrangian Method
We provide a direct comparison between our PINN-MEP approach and the method of Du et al. in terms of computational efficiency, as evaluated by the number of force field/energy function calls, and the maximum energies observed in the paths, corresponding to the peak of the potential barrier. To provide a rigorous comparison, we conducted benchmarks using experimental conditions identical to those reported by Du et al. for alanine dipeptide: AMBER14 force field, 300K temperature. Table 1 presents these results alongside those reported in Du et al. (2024) and their recent follow-up work Lee et al. (2025). The only exception is that we do not evaluate trajectory probability, as the assumption of optimization over a non-physical path length makes the calculation of this quantity impossible.
@@ Appendix B Comparison with other related Methods
While our primary benchmark comparison focuses on the Doob’s Lagrangian method due to similarity in task and system setup, examining how our approach relates to other methods for transition path discovery/MEP generation is also informative. Here we present results from two related approaches in the literature, though we emphasize that direct numerical comparisons should be treated with caution due to differences in force fields, optimization objectives, and units used.
Method | # Evaluations | Max Energy (eV) |
NEB | 8063 | 11.0014 |
@@
Method | CVs | # Evaluations | Runtime / Path |
@@
Table 2 shows results from Ramakrishnan et al. (2025) comparing Nudged Elastic Band (NEB) with their Implicit Neural Representation (INR) approach for MEP discovery. Both our PINN-MEP approach and their INR method utilize lightweight, preexisting molecular dynamics force fields, though we employ different such force fields. Their approach requires a more complex loss function that evaluates both energy and forces along with path parameterization and its gradient (path velocity), but needs fewer force field evaluations than our method. It is important to note that the "Max Energy" reported in Table 2 is measured in electron volts (eV) and appears to be adjusted for ground state energy reference, which makes direct energy comparisons with our method challenging. The difference in force fields used further complicates any direct energy value comparisons between methods.
In contrast, Table 3 presents results from Raja et al. (2025), who take a fundamentally different approach by re-purposing pre-trained generative models to perform sampling of transition paths. A key advantage of their method is that the generative model provides significantly better initial guesses for transition paths based on learned molecular distributions. Their approach requires evaluating both the pretrained model (as a surrogate force field) and its gradient for their loss function, which is more computationally intense due to the high number of parameters of the pretrained model as compared to a physics-derived force field.
These tables corroborate our findings in the Doob’s Lagrangian comparison: methods optimizing for a single MEP consistently require orders of magnitude fewer force field evaluations than approaches that characterize the full transition path ensemble.
@@ B.1 Ablation Studies
This section presents a series of ablation studies examining the impact of various components of our PINN-MEP approach. These experiments help validate our design choices and provide insights into the importance of different architectural and training elements. We conduct these ablation studies on the AIB9 system because Alanine dipeptide, as a benchmark system, is too simple to provide meaningful insights into the performance of our method. For each ablation study, we report the lowest maximum energy of the path among all training runs, the mean and standard deviation of the energy, and the mean number of iterations to reach the lowest maximum energy. We run each setting in the ablations 15 times for a maximum of 25,000 iterations. We use a batch size of 1, where each batch contains 32 values of the progress coordinate for most cases. The exception is when we evaluate the chain-of-states-like regularization, which minimizes the difference between network outputs for adjacent values of ; in this case, the effective batch size is 32, since there are no interactions between different network outputs otherwise.
Due to the nature of physical force fields, slight deviations in the path can lead to large changes in energy, leading to similar settings having very different energies. This can limit the ability to compare the performance of different settings. However, we find that the results are still directionally informative.
B.1.1 General Settings for Ablation Studies
All ablation studies are conducted using the following base configuration, with each study varying only specific parameters while keeping all others fixed:
Molecular System Parameters:
Molecular system: AIB9 peptide with 129 atoms
Force field: AMBER ff15ipq-m for protein mimetics
Cutoff method: No cutoff applied
Simulation box size: 10.0 nm
Neighbor list cutoff distance: 4.0 nm
Training Configuration:
Number of iterations: 25,000 per run
Number of independent runs: 15 per parameter setting
Evaluation frequency: Every 100 training steps
Optimizer: Adam with optax default parameters
Learning rate:
Path Sampling Parameters:
Batch size: 1
Frames per batch: 32 points sampled along the progress coordinate
Spring constant: (no chain-of-states regularization)
Interpolation method: Cubic Hermite for blending functions
Network Architecture (Base Configuration):
Network depth: 6 hidden layers
Width factor: 4 (network width scales as number of atoms)
Embedding dimension: Automatically set to
Time embedding: Sinusoidal with maximum period of
Activation function: ReLU throughout the network
Weight initialization: Glorot uniform for hidden layers, zero initialization for final layer
Bias initialization: Default for hidden layers, zero initialization for final layer
B.1.2 Embedding Strategy
We compared sinusoidal time embeddings with different maximum periods against simple linear projections of the progress coordinate. Table 4 shows that linear embedding (None) performed best, achieving the lowest mean energy, though sinusoidal embeddings with period also achieved competitive results. We find these results at odds with the need for sinusoidal embeddings in the larger BPTI system. Perhaps the AIB9 system, which is characterized by one large general conformational rearrangement primarily driven by main-chain dihedral angles, does not need the high frequency features that sinusoidal embeddings provide, but which might be essential for other systems, due to the well-known bias of MLPs toward modeling low frequencies in data.
Embedding | Max Period | MeanMax Energy () | MinMax Energy | Mean Iterations |
Linear (None) | – | 591.35 963.72 | -337.36 | 200873 |
Sinusoidal | 1326.13 2339.35 | -337.36 | 210907 | |
Sinusoidal | 1528.18 2429.05 | -129.90 | 226290 | |
Sinusoidal | 1888.66 3604.47 | -337.36 | 187493 | |
Sinusoidal | 2453.53 3507.47 | -337.36 | 217053 |
B.1.3 Spring Regularization
We tested the effect of chain-of-states-like spring constant regularization, as shown in Table 5. Interestingly, we found that small amounts of spring regularization ( or ) provided the best results, with achieving the lowest mean energy overall. This suggests that a minimal amount of explicit path continuity enforcement can be beneficial when combined with our neural representation, though this is not a strict requirement for the method to work. However, adding a spring constant reintroduces intra-batch dependencies, which means that batch sizes should be of a minimum size for this approach to work effectively. With low batch sizes, this could lead to the corner cutting effects commonly observed in traditional chain-of-states methods, which might make scaling the addition of spring regularization difficult for larger systems. Since we are able to reach the same minimum energy without spring regularization, we recommend against its use, especially when scaling to larger systems where lower batch sizes are preferred.
Spring Constant | MeanMax Energy () | MinMax Energy | Mean Iterations |
-178.22 260.09 | -337.36 | 153533 | |
-155.41 286.63 | -337.36 | 101600 | |
803.61 586.95 | 205.74 | 159033 | |
0.0 | 1460.41 2486.05 | -337.36 | 209133 |
10528.4 7394.89 | 5065.46 | 155400 | |
48639.7 12111.7 | 33143.6 | 107267 |
B.1.4 Network Architecture
@@
We investigated the effect of network depth and width on transition path quality. Table 6 shows that an 8-layer network with a width factor of 4 achieved the best mean energy performance, followed closely by a 6-layer network with width factor 6.
@@
Depth | Width Factor | MeanMax Energy () | MinMax Energy | Mean Iterations |
8 | 4 | 25.02 467.23 | -337.36 | 159333 |
6 | 6 | 72.42 458.72 | -337.36 | 210267 |
8 | 6 | 165.91 1021.79 | -337.36 | 219433 |
8 | 2 | 1434.97 3206.06 | -337.36 | 178167 |
6 | 4 | 1460.34 2486.07 | -337.36 | 209067 |
6 | 2 | 3621.8 3198.25 | 217.83 | 231200 |
4 | 6 | 4242.74 3568.51 | 140.81 | 231967 |
4 | 4 | 5130.75 5758.01 | -328.49 | 228133 |
4 | 2 | 16200.5 8032.18 | 159.21 | 210367 |
B.1.5 Interpolation Type
We examined different interpolation methods for the blending functions that combine the neural network output with the endpoint constraints. Table 7 compares linear, cosine, quintic, and cubic Hermite interpolation. Cosine interpolation performed best in terms of mean energy, though all methods were able to achieve the minimum energy of -337.36 in at least one run, except for quintic. The cubic Hermite interpolation showed higher variance in results.
The specific formulas for each interpolation type are as follows:
Linear interpolation:
Cubic Hermite interpolation:
Quintic interpolation:
Cosine interpolation:
where and are the basis functions that ensure endpoint constraints and , and is the blending function that controls the neural network’s contribution along the path.
Interpolation Type | MeanMax Energy () | MinMax Energy | Mean Iterations |
Cosine | 442.51 769.51 | -337.36 | 198720 |
Linear | 602.61 1057.13 | -337.36 | 219147 |
Quintic | 967.16 1288.57 | -281.14 | 210227 |
Cubic Hermite | 1325.49 2339.75 | -337.36 | 212807 |
B.1.6 AIB9 Network Architecture
Below are the main details of the AIB9 MEP training runs, as shown in the main paper.